# Prediction Intervals for Interim Events in Randomized Clinical Trials with Time-to-Event Endpoints

### v1

@@ 3.2 Proposed Sampling Scheme
The original DB–MC method requires generating many replicates of the multiple-cohort dataset and, for each replicate, computing the vector of conditional probabilities $\tilde{\bm{\pi}}$. Translating this step to interim monitoring—where each patient can be viewed as a cohort of size one—requires additional care. We propose a parametric bootstrap sampling scheme to obtain the bootstrap estimator if the conditional CMF of $Y$. The proposed scheme is exposed in the Appendix, Section 7.5.
Some remarks are worth noting regarding the nature of the inference considered in this particular contest. The proposed sampling scheme conditions on the event and censoring indicators, i.e, they are kept fixed across bootstrap replicates. If censoring indicators were different as in classic survival bootstrap [23] [36] [1] [30] [11], the risk set would vary across replicates and the Monte Carlo average in (3) would mix conditional CMFs with different supports. This would target an unconditional CMF, whereas in a within-sample prediction problem it is required the CMF of $Y$ conditional on the observed risk set that defines its support. Conditioning on the indicators also brings information on the observed times. Moreover, each $\pi_{j}$ is conditional on $\tau_{j}$; hence, the bootstrap must preserve this conditioning in order to recover the sampling distribution of $\tilde{\boldsymbol{\pi}}$. Prior observing the interim data, $(T_{j}^{\mathrm{obs}},\delta_{j},\varepsilon_{j},\gamma_{j})_{j=1}^{n}$ are random variables with PDF $\mathrm{Pr}_{\bm{\vartheta}}(T_{j}^{\mathrm{obs}},\delta_{j},\varepsilon_{j},\gamma_{j}\mid\bm{z}_{j})$. Once observing the indicators values and fixing the entry dates, we have
$\mathrm{Pr}\bigl(T_{j}^{\mathrm{obs}}=t\mid\delta_{j}=1,\bm{z}_{j}\bigr)$ $=\frac{f_{\bm{\theta}}(t\mid\bm{z}_{j})}{F_{\bm{\theta}}(\tau_{j},\bm{z}_{j}\bigr)}\,$, (9)
@@
It follows that event times and loss-to-follow-up times are sampled from their respective truncated PDFs in (9), whereas interim censoring times are kept fixed at their realized values, thus ignoring how they are distirbuted. Sampling event times and loss-to-follow-up times from (9) can be carried out straightforwardly via inverse-transform. This approach is flexible and accommodates a wide range of lifetime distributions, requiring only evaluation of the corresponding truncated PDF and its inverse [9] [20]. See Appendix, Section 7.5 for further details. In each bootstrap replicate, we obtain $\widehat{\bm{\theta}}$ and $\widehat{\bm{\psi}}$ by maximizing the likelihood in (8) under the same assumptions, and then compute $\pi_{j}$ accordingly.
@@ 4 Simulation Study
We conducted a main simulation study, namely $\mathcal{S}_{1}$, to globally assess the operating characteristics of the proposed method under various conditions of interest in the interim analysis contest. We also conducted an additional simulation study $\mathcal{S}_{2}$ inspired by the real case in which we focus on the impact of potential model over-parametrization on the operating characteristics of the proposed method. Our simulations not only assess the impact of the proposed extensions, but also investigate the behavior of the original DB-MC method. Without loss of generality, we assume that data are measured in years in each data-generating process.
@@ 4.1 Data generating process
In $\mathcal{S}_{1}$ data were generated from Weibull proportional hazards model with uniform entry dates over three years of recruitment period. In addition, we introduce an exponential loss-to-follow-up mechanism that is correlated with entry time, so that patients enrolled earlier are more likely to be lost to follow-up than those enrolled later. To induce this dependence while keeping the entry-time and loss-to-follow-up marginals PDFs parameterized separately, we use NORTA sampling (NORmal To Anything) [14] and calibrate the target correlation via a bisection procedure [5] with tolerance 0.001 and a maximum of 100 iterations with sample size 20000. The Weibull shape parameter was fixed at 0.6 (decreasing hazard). The baseline scale parameter was varied so that the proportion of interim censored observations matched the target censoring proportion under a specific value of the hazard ratio (see next Section), using a bisection procedure [5] with starting value 0, tolerance 0.001, and a maximum of 300 iterations. We considered one independent subject-level baseline binary covariate $Z$ generated from a Bernoulli distribution with parameter 0.5, i.e., we considered only the treatment covariate with equal randomized patient between two arms. The Poisson–binomial distribution was approximated in each bootstrap sample by a Poisson distribution, as Hong [39] showed this approximation to be highly accurate for a moderately large number of Bernoulli trials (e.g. >200) while being computationally more efficient than exact methods.
In $\mathcal{S}_{2}$, the data-generating process is the same as in $\mathcal{S}_{1}$. For this study, we imposed only administrative censoring at the interim cutoff. Full details of the generating mechanisms and the resulting parameter values are reported in the Appendix, Section 7.6.
@@ 4.2 Factors in the simulations
In $\mathcal{S}_{1}$, the performance of the proposed method is evaluated over 72 scenarios defined by a combination of the following factors: the correlation between the lost to follow-up distribution and the entry dates $\rho=\{0.1,0.5\}$, the follow-up maturity at the time of prediction, $t_{c}=\{1,2,4\}$, expressed as the time from accrual end to the interim look; the ratio between the baseline Weibull scale and the loss to follow-up rate, $k=\{0.1,0.2,0.3\}$; the strength of the treatment effect, expressed by the hazard ratio, $HR=\{0.2,0.8\}$; and the future prediction horizon, $\Delta t=\{1,4\}$. The proportion of interim censored observations is fixed at 0.5. Here, we evaluate prediction intervals under correct model specification.
In $\mathcal{S}_{2}$, the methodology is evaluated over 24 scenarios defined by a combination of the following factors: the proportion of interim censored observations at the interim, $0.2=\{0.2,0.8\}$ and $t_{c}$, $HR$, $\Delta t$ using the same values defined in $\mathcal{S}_{1}$. We evaluated prediction intervals obtained from the Weibull model, Royston–Parmar models with 1, 2, and 3 knots placed at the 25th, 50th, and 75th percentiles of the uncensored event times and the Generalized Gamma model [67], which encompasses Weibull and log-normal as special cases [59] [19]. Following Mori et al. [56], we decided to include Royston–Parmar models [62] as a flexible extension of standard parametric regression models. These models generalize the linear relationship between a transformed survival function, i.e. a link function $g(S(t))$ and log time $y=\log(t)$ on a chosen covariates effect scale (e.g. proportional hazard scale) by modeling this relationship with restricted cubic splines. In this sense, they extend the Weibull proportional hazards model, the log-logistic proportional odds model, and the log-normal linear probit model. Despite their added flexibility, Royston–Parmar models preserve an explicit baseline specification, allowing extrapolation beyond the last observed event. Flexibility is controlled by the number of internal knots placed between boundary knots [62].
@@ 4.3 Performance metrics
The prediction intervals obtained within the DB-MC framework enjoy the classical frequentist interpretation: they correspond to the long‐run proportion of times that $Y$ falls in a random region $\mathrm{PI}_{1-\alpha}(\mathscr{D})$, i.e.the prediction interval, under repeated sampling from the same data‐generating mechanism [8][50]. Coverage probability (CP) is a useful performance metric to evaluate an interval estimator. Since $Y$ is also random, there are two natural notions of CP associated with a prediction interval procedure [10] [54]. First, for the observed data $\boldsymbol{D}$ there is the coverage probability of the realized prediction interval, i.e., the conditional coverage probability (CCP)
@@
where $\bm{D}^{*}_{1},\ldots\bm{D}^{*}_{N}$ are simulated data from a given model. In our simulation studies, $\bm{\vartheta}=(\bm{\theta},\bm{\psi},\bm{\phi})$ and $\varphi(\boldsymbol{\vartheta})=\tilde{\bm{\pi}}$. The UCP of the prediction intervals is estimated under the factors described in Section 4.2 using the following steps:
@@
    For each $\boldsymbol{D}^{*}_{k}$, we evaluate the CCP using the conditional distribution of $Y$ evaluated at the true $\tilde{\bm{\pi}}$, i.e. the $\tilde{\bm{\pi}}$ evaluated on the simulated dataset $\boldsymbol{D}^{*}_{k}$ at the design value of $\bm{\vartheta}$.
@@
[…]
@@ 4.4 Results
Figure 1: Study $\mathcal{S}_{1}$. Performance metrics of the 95% two-sided prediction intervals under different factors. $\Delta t$ denotes the prediction horizon, $t_{c}$ the time of interim from accrual end, $\rho$ the correlation between lost to follow-up times and entry dates, HRHR is the hazard ratio and $k$ the ratio between the baseline scale parameter and the loss to follow-up rate.
@@ 5 Case study in Pediatric ALL
For the application of our method, two-sided prediction intervals at the 0.95 confidence level were obtained for the the number of additional EFS events from June 1, 2017 (accrual closure) and updated every six months until June 1, 2021, when the final analysis was scheduled. The full dataset was used to evaluate if the observed events are included by the prediction intervals. EFS is defined as the time from random assignment to the date of the first event (resistance to treatment, relapse, death, or second malignant neoplasm, whichever occurred first) or to the last follow-up. Since no lost to follow-up occurred at the time of the final analysis, all patients who had not experienced an EFS event by June 1, 2017 were considered administratively censored at that date. The presence of interim censoring only leads to a simplified formulation of (6) and the sampling scheme, as detailed in the Appendix, Section 7.3. On June 1 2017, 164 events had occurred among 809 enrolled patients. By the final analysis, 62 further events had been observed, for a total of 226 events over the entire trial.
@@
Figure 3 shows the relationship between different transformations of the survival function and natural logarithm of time: straight parallel lines suggest correct specification of both baseline survival and covariates effect scale. These plots showed that there is no linear relationships between the transformed survival and the logarithm of time, suggesting that Royston-Parmar models may be adequate. In addition, PH assumption are untenable. The smoothed hazard plot (Figure 8, Appendix) indicates decreasing hazards converging later in the follow-up.
@@
Figure 3: Different survival function transformations vs. log-time by treatment arm, 1 = Active treatment, 0 = standard of care. Panel A shows the log cumulative hazard of EFS $\mathrm{log}(-\mathrm{log}(S(t)))$; Panel B shows the cumulative probit of EFS $\Phi^{-1}(S(t))$, where $\Phi^{-1}$ is che CDF of the standard normal distirbution; Panel C shows the cumulative logit of EFS $\mathrm{log}(1-S(t)/S(t))$
Figure 4: Two-sided 95% prediction intervals for the additional number of EFS events from June 1st 2017 to June 1st 2021 . Black diamonds indicate the observed counts. Panel A shows intervals under the Weibull model and its flexible specifications; Panel B under the log-normal model and its flexible specifications; Panel C under the log-logistic model and its flexible specifications. RP = Royston–Parmar; $k$ denotes the number of knots.
@@ 6 Discussion and Conclusions
The Poisson–binomial representation of the future event count provides a flexible way to model heterogeneous, subject-specific conditional probabilities of experiencing the event within the future time window, driven by staggered entry and, more generally, by covariate profiles. This formulation also naturally allows different parametric models for event times and censoring times. Embedding such models within the sampling scheme proposed is straightforward since it requires simulating event and lost to follow-up times from the chosen truncated parametric distributions. This yields substantial flexibility, leveraging a broad class of parametric survival models available in the literature [21] [51]. In tjos work we assumed a common dropout mechanism across covariate profiles, but the approach readily extends to settings where dropout also varies by covariates (e.g., treatment arm). The key requirement is that, conditional on the covariates, the event time and the loss-to-follow-up time are independent and their distributions share no parameters.
We examined the operating characteristics of the proposed method across a range of factors of interested in interim monitoring settings and with a case study on pediatric ALL. Follow-up maturity and prediction horizon emerged as key determinants of interval performance: more mature follow-up and shorter horizons led to better coverage, whereas longer horizons reduced coverage. Loss to follow-up produced modest reduction in coverage, which became more pronounced for longer prediction horizons. This reflects an information trade-off: while additional loss to follow-up can inform the censoring model, it simultadditional number of eventsously reduces information for estimating the event-time distribution because fewer events are observed.
@@
Further research could address the limitations underlying our approach and connect it to a set of active areas of research. First, we assume that the data at the interim analysis is free of reporting delays. In multicenter trials, however, delays in data entry and event reporting are common and can materially affect interim monitoring and prediction [73] [75] [4]. Extensions that explicitly account for delayed reporting would therefore be valuable. Second, we focus on settings in which accrual is complete at the prediction time. In many trials, interim analyses occur while recruitment is ongoing, so a practically relevant extension is joint prediction of both future accrual and future events as often considered in literature on time to endpoint maturation. Third, we assume access to unblinded, patient-level information. In practice, interim analyses are frequently conducted under blinding constraints to preserve trial integrity, limiting the availability of individual-level covariates or even treatment indicators. Related work on time to endpoint maturation prediction under blinded data [22] [26] [84] suggests promising directions to adapt our framework to blinded settings. Regarding the survival model, we restrict attention to regression-based parametric specifications without explicitly modeling potential violations of the covariate effect scale (e.g., departures from proportional hazards, proportional odds, or accelerated failure time assumptions). While in the motivating application ignoring such violations did not lead to marked differences in prediction intervals, the generalizability of this observation warrants further investigation. Lastly, our work may benefit from the extensive literature on parametric survival extrapolation from clinical trials for economic evaluation [49, 64, 29, 18, 48, 68, 40, 58, 15], where model choice and extrapolative performance are central concerns. Leveraging insights from this literature could be a promising avenue for future work.

### v2

@@ 3.2 Proposed Sampling Scheme
The original DB–MC method requires generating many replicates of the multiple-cohort dataset and, for each replicate, computing the vector of conditional probabilities $\bm{\pi}$ at a future prediction horizon $\Delta t$. Translating this step to interim monitoring—where each patient can be viewed as a cohort of size one—requires additional care. We propose a parametric bootstrap sampling scheme to obtain the bootstrap estimator if the conditional CMF of $Y$. The proposed scheme is exposed in the Appendix, Section 7.5.
Some remarks are worth noting regarding the nature of the inference considered in this particular contest. The proposed sampling scheme conditions on the event and censoring indicators, i.e, they are kept fixed across bootstrap replicates. If censoring indicators were different as in classic survival bootstrap [23, 36, 1, 30, 11], the risk set would vary across replicates and the Monte Carlo average in (5) would mix conditional CMFs with different supports. This would target an unconditional CMF, whereas in a within-sample prediction problem it is required the CMF of $Y$ conditional on the observed risk set that defines its support. Conditioning on the indicators also brings information on the observed times. Moreover, each $\pi_{j}$ is conditional on $\tau_{j}$. Hence, the bootstrap must preserve this conditioning in order to recover the sampling distribution of $\boldsymbol{\pi}$ at a future prediction horizon $\Delta t$. Prior observing the interim data $\bm{D}$, $(T_{j}^{\mathrm{obs}},\delta_{j},\varepsilon_{j},\gamma_{j})_{j=1}^{n}$ are random variables with PDF $\mathrm{Pr}_{\bm{\vartheta}}(T_{j}^{\mathrm{obs}},\delta_{j},\varepsilon_{j},\gamma_{j}\mid\bm{z}_{j})$. Once observing the indicators values and fixing the entry dates, we have
$\mathrm{Pr}\bigl(T_{j}^{\mathrm{obs}}=t\mid\delta_{j}=1,\bm{z}_{j}\bigr)$ $=\frac{f_{\bm{\theta}}(t\mid\bm{z}_{j})}{F_{\bm{\theta}}(\tau_{j}\mid\bm{z}_{j}\bigr)}\,$, (11)
@@
It follows that event times and loss-to-follow-up times are sampled from their respective truncated PDFs in (11), whereas interim censoring times are kept fixed at their realized values, thus ignoring how they are distirbuted. Sampling event times and loss-to-follow-up times from (11) can be carried out straightforwardly via inverse-transform. This approach is flexible and accommodates a wide range of lifetime distributions, requiring only evaluation of the corresponding truncated PDF and its inverse [9] [20].See Appendix, Section 7.5 for further details. In each bootstrap replicate, we obtain $\widehat{\bm{\theta}}$ and $\widehat{\bm{\psi}}$ by maximizing the likelihood in (10) under the same assumptions, and then compute $\pi_{j}$ at a future prediction horizon $\Delta t$ accordingly.
@@ 4 Simulation Study
We conducted two comprehensive simulation studies to globally assess the operating characteristics of the proposed method under various conditions of interest in the interim analysis contest. Our simulations not only assess the impact of the proposed extensions, but also investigate the behavior of the original DB-MC method.
@@ 4.1 Data generating process
We considered two simulation studies with different data generating processes and experimental conditions, namely $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$. In $\mathcal{S}_{1}$ data were generated from Weibull proportional hazards model with uniform entry dates over three years of recruitment period. In addition, we introduce an exponential loss-to-follow-up mechanism that is correlated with entry time, so that patients enrolled earlier are more likely to be loss to follow-up than those enrolled later. To induce this dependence while keeping the entry-time and loss-to-follow-up marginals PDFs parameterized separately, we use NORTA sampling (NORmal To Anything) [14] and calibrate the target correlation via a bisection procedure [5] with tolerance 0.001 and a maximum of 100 iterations with sample size 20000. The Weibull shape parameter was fixed at 0.6 (decreasing hazard). The baseline scale parameter was varied so that the proportion of interim censored observations matched the target censoring proportion under a specific value of the hazard ratio (see next Section), using also a bisection procedure [5] with starting value 0, tolerance 0.001, and a maximum of 300 iterations. We considered treatment as covariate, assuming a 1:1 randomization ratio generated from a Bernoulli distribution with parameter 0.5. The Poisson–binomial distribution was approximated in each bootstrap sample by a Poisson distribution, as Hong [39] showed this approximation to be highly accurate for a moderately large number of Bernoulli trials (e.g. >200) while being computationally more efficient than exact methods.
In $\mathcal{S}_{2}$, the data-generating process is similar $\mathcal{S}_{1}$ but we imposed only administrative censoring at the interim cutoff. Full details of the generating mechanisms and the resulting parameter values are reported in the Appendix, Section 7.6.
@@ 4.2 Factors in the simulations
In $\mathcal{S}_{1}$, the performance of the proposed method is evaluated over 72 scenarios defined by a combination of the following factors: the correlation between the loss to follow-up distribution and the entry dates $\rho=\{0.1,0.5\}$, the follow-up maturity at the time of prediction, $t_{c}=\{1,2,4\}$, expressed as the time from accrual end to the interim look; the ratio between the baseline Weibull scale and the loss to follow-up rate, $k=\{0.1,0.2,0.3\}$; the strength of the treatment effect, expressed by the hazard ratio, $HR=\{0.2,0.8\}$; and the future prediction horizon, $\Delta t=\{1,4\}$. The proportion of interim censored observations is fixed at 0.5. Here, we evaluate prediction intervals under correct model specification.
In $\mathcal{S}_{2}$, the methodology is evaluated over 24 scenarios defined by a combination of the following factors: the proportion of interim censored observations, $0.2=\{0.2,0.8\}$ and $t_{c}$, $HR$, $\Delta t$ using the same values defined in $\mathcal{S}_{1}$. We evaluated prediction intervals obtained from the Weibull model, Royston–Parmar models with 1, 2, and 3 knots placed at the 25th, 50th, and 75th percentiles of the uncensored event times and the Generalized Gamma model [67], which encompasses Weibull and log-normal as special cases [59] [19]. Following Mori et al. [56], we decided to include Royston–Parmar models [62] as a flexible extension of standard parametric regression models. These models generalize the linear relationship between a transformed survival function, i.e. a link function $g(S(t))$ and log time $y=\log(t)$ on a chosen covariates effect scale (e.g. proportional hazard scale) by modeling this relationship with restricted cubic splines. In this sense, they extend the Weibull proportional hazards model, the log-logistic proportional odds model, and the log-normal linear probit model. Despite their added flexibility, Royston–Parmar models preserve an explicit baseline specification, allowing extrapolation beyond the last observed event. Flexibility is controlled by the number of internal knots placed between boundary knots [62].
@@ 4.3 Performance metrics
The prediction intervals obtained within the DB-MC framework enjoy the classical frequentist interpretation: they correspond to the long‐run proportion of times that $Y$ falls in a random region $\mathrm{PI}_{1-\alpha}(\mathscr{D})$, i.e.the prediction interval, under repeated sampling from the same data‐generating mechanism [8, 50]. Coverage probability (CP) is a useful performance metric to evaluate an interval estimator. Since $Y$ is also random, there are two natural notions of CP associated with a prediction interval procedure [10, 54]. First, for the observed data $\boldsymbol{D}$ there is the coverage probability of the realized prediction interval, i.e., the conditional coverage probability (CCP)
@@
where $\bm{D}^{*}_{1},\ldots\bm{D}^{*}_{N}$ are simulated data from a given model. In our simulation studies, $\bm{\vartheta}=(\bm{\theta},\bm{\psi},\bm{\phi})$ and $\varphi(\boldsymbol{\vartheta})={\bm{\pi}}$. The UCP of the prediction intervals is estimated under the factors described in Section 4.2 using the following steps:
@@
    For each $\boldsymbol{D}^{*}_{k}$, we evaluate the CCP using the conditional distribution of $Y$ evaluated at the true $\bm{\pi}$}, i.e. the $\bm{\pi}$} evaluated on the simulated dataset $\boldsymbol{D}^{*}_{k}$ at the design value of $\bm{\vartheta}$.
@@
$\begin{gathered}\frac{y_{\alpha}\bigl(\boldsymbol{D}_{k}^{*}\bigr)-y_{\alpha}\bigl(\boldsymbol{D}_{k}^{*},{\bm{\pi}})}{y_{\alpha}\bigl(\boldsymbol{D}_{k}^{*},{\bm{\pi}}\bigr)}\,,\\ \\[4.0pt] \frac{\mathrm{PI}_{1-\alpha}(\boldsymbol{D}_{k}^{*})}{\mathrm{PI}_{1-\alpha}(\boldsymbol{D}_{k}^{*},{\bm{\pi}})}\,.\end{gathered}$ (15)
where $y_{\alpha}\bigl(\boldsymbol{D}_{k}^{*},\bm{\pi})$ and $\mathrm{PI}_{1-\alpha}(\boldsymbol{D}_{k}^{*},{\bm{\pi}})$ are the $\alpha$ quantile and the prediction interval obtained on the simulated dataset $\bm{D}_{k}^{*}$ at the true value of $\bm{\pi}$}.
@@ 4.4 Results
Figure 1: Study $\mathcal{S}_{1}$. Performance metrics of the 95% two-sided prediction intervals under different factors. $\Delta t$ denotes the prediction horizon, $t_{c}$ the time of interim from accrual end, $\rho$ the correlation between loss to follow-up times and entry dates, HRHR is the hazard ratio and $k$ the ratio between the baseline scale parameter and the loss to follow-up rate.
@@ 5 Case study in Pediatric ALL
For the application of our method, two-sided prediction intervals at the 0.95 confidence level were obtained for the the number of additional EFS events from June 1, 2017 (accrual closure) and updated every six months until June 1, 2021, when the final analysis was scheduled. The full dataset was used to evaluate if the observed events are included by the prediction intervals. EFS is defined as the time from random assignment to the date of the first event (resistance to treatment, relapse, death, or second malignant neoplasm, whichever occurred first) or to the last follow-up. Since no loss to follow-up occurred at the time of the final analysis, all patients who had not experienced an EFS event by June 1, 2017 were considered administratively censored at that date. The presence of interim censoring only leads to a simplified formulation of (8) and the sampling scheme, as detailed in the Appendix, Section 7.3. On June 1 2017, 164 events had occurred among 809 enrolled patients. By the final analysis, 62 further events had been observed, for a total of 226 events over the entire trial.
@@
Figure 3 shows the relationship between different transformations of the survival function and natural logarithm of time: straight parallel lines suggest correct specification of both baseline survival and covariates effect scale. These plots showed that there is no linear relationships between the transformed survival and the logarithm of time, suggesting that Royston-Parmar models may be adequate. In addition, PH assumptions are untenable. The smoothed hazard plot (Figure 8, Appendix) indicates decreasing hazards converging later in the follow-up.
@@
Figure 3: Different survival function transformations vs. log-time by treatment arm, 1 = Active treatment, 0 = standard of care. Panel A shows the log cumulative hazard of EFS $\mathrm{log}(-\mathrm{log}(S(t)))$; Panel B shows the cumulative probit of EFS $\Phi^{-1}(S(t))$, where $\Phi^{-1}$ is CDF of the standard normal distribution; Panel C shows the cumulative logit of EFS $\mathrm{log}(1-S(t)/S(t))$.
Figure 4: Two-sided 95% prediction intervals for the additional number of EFS events from June 1st 2017 to June 1st 2021 . Black diamonds indicate the observed counts. Panel A shows intervals under the Weibull model and its flexible specifications; Panel B shows under the log-normal model and its flexible specifications; Panel C shows under the log-logistic model and its flexible specifications. RP = Royston–Parmar; $k$ denotes the number of knots.
@@ 6 Discussion and Conclusions
The Poisson–binomial representation of the future event count provides a flexible way to model heterogeneous, subject-specific conditional probabilities of experiencing the event within the future time window, driven by staggered entry and, more generally, by covariate profiles. This formulation also naturally allows different parametric models for event times and censoring times. Embedding such models within the sampling scheme proposed is straightforward since it requires simulating event and loss to follow-up times from the chosen truncated parametric distributions. This yields substantial flexibility, leveraging a broad class of parametric survival models available in the literature [21, 51]. In this work we assumed a common dropout mechanism across covariate profiles, but the approach readily extends to settings where dropout also varies by covariates (e.g., treatment arm). The key requirement is that, conditional on the covariates, the event time and the loss-to-follow-up time are independent and their distributions share no parameters.
We examined the operating characteristics of the proposed method across a range of factors of interested in interim monitoring settings and with a case study on pediatric ALL. Follow-up maturity and prediction horizon emerged as key determinants of interval performance: more mature follow-up and shorter horizons led to better coverage, whereas longer horizons reduced coverage. Loss to follow-up produced modest reduction in coverage, which became more pronounced for longer prediction horizons. This reflects an information trade-off: while additional loss to follow-up can inform the censoring model, it simultaneously reduces the information for estimating the event-time distribution because fewer events are observed.
@@
Further research could address the limitations underlying our approach and connect it to a set of active areas of research. First, we assume that the data at the interim analysis is free of reporting delays. In multicenter trials, however, delays in data entry and event reporting are common and can materially affect interim monitoring and prediction [73, 75, 4]. Extensions that explicitly account for delayed reporting would therefore be valuable. Second, we focus on settings in which accrual is complete at the prediction time. In many trials, interim analyses occur while recruitment is ongoing, so a practically relevant extension is joint prediction of both future accrual and future events as often considered in literature on time to endpoint maturation. Third, we assume access to unblinded, patient-level information. In practice, interim analyses are frequently conducted under blinding constraints to preserve trial integrity, limiting the availability of individual-level covariates or even treatment indicators. Related work on time to endpoint maturation prediction under blinded data [22, 26, 84] suggests promising directions to adapt our framework to blinded settings. Regarding the survival model, we restrict attention to regression-based parametric specifications without explicitly modeling potential violations of the covariates effect scale (e.g., departures from proportional hazards, proportional odds, or accelerated failure time assumptions). While in the motivating application ignoring such violations did not lead to marked differences in prediction intervals, the generalizability of this observation warrants further investigation. Lastly, our work may benefit from the extensive literature on parametric survival extrapolation from clinical trials for economic evaluation [49, 64, 29, 18, 48, 68, 40, 58, 15], where model choice and extrapolative performance are central concerns. Leveraging insights from this literature could be a promising avenue for future work.